OpenStudy (anonymous):

Solve. 8=2^(x+4) Help :(

OpenStudy (turingtest):

eight is two to what power?

OpenStudy (anonymous):

to the third power

OpenStudy (turingtest):

good, so write that

OpenStudy (turingtest):

\[2^3=2^{x+4}\]

OpenStudy (anonymous):

like 2^3=2^(x+4)?

OpenStudy (turingtest):

yep now, for both sides to be equal the exponents must be equal make sense?

OpenStudy (anonymous):

yes that makes sense. but how can you subtract the exponents

OpenStudy (turingtest):

just write out the exponents and take out the common base of 2

OpenStudy (turingtest):

what is the exponent on the left? what is the exponent on the right? set them equal and solve for x

OpenStudy (anonymous):

left exponent is 3, right is x+4 OHH so 3=x+4 -4 x=-1?

OpenStudy (turingtest):

yes :)

OpenStudy (anonymous):

thank you!!!:)

OpenStudy (turingtest):

very welcome!

OpenStudy (kainui):

This made me smile hah.

hartnn (hartnn):

5 medals !! are u kidding me !

OpenStudy (anonymous):

It makes sense to me.. Just look for the way he taught him.. One from my side too..